Q:-If p=λ(u×v)+μ(v×w)+κ(w×u) and [uvw]=1/5,then find λ+μ+κ.(here p,u,v,w are vectors and λ,μ,κ are scalars)


a. 5


b. 10


c. 15


d. none

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Tapasranjan Das

p=λ(u×v)+μ(v×w)+κ(w×u)

then

u.P =μ u.(vXw)

       =μ [u v w]

        =μ/5

similerly

  v.P = κ/5

and  w.P = λ/5

 add these value   (λ+μ+κ)/5 =u.P + v.P + w.P

                                 λ+μ+κ   =5(u+v+w).P

   clearly different combination of u,v,w final result will change .

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin